Problem B.1: Temperature of Earth (6 Points)

Our Sun shines bright with a luminosity of 3.828 x 10 Watt. Her energy is responsible for many processes and the habitable temperatures on the Earth that make our life possible.

(a) Calculate the amount of energy arriving on the Earth in a single day.

litres of heating oil (energy density: 37.3 x 10 J/litre) is this equivalent?

(b) To how many (c) The Earth reflects 30% of this energy: Determine the temperature on Earth's surface.

(d) What other factors should be considered to get an even more precise temperature estimate?

Note: The Earth's radius is

6370 km, the Sun's radius is 696 x 10 km; 1 AU is 1.495 x 10 km.
1
Expert's answer
2020-06-22T11:18:38-0400

(a) Calculate the amount of energy arriving on 1 square meter of the Earth:


"E_0=\frac{L}{A}=\frac{L_S}{4\pi R_S^2}=\frac{3.828⋅10^{26}}{4\pi (1.495⋅10 ^{11})^2}=1360\text{ W/m}^2."


This power comes to the earth's cross-section. Multiply by time and get energy:


"E=E_0\pi R_E^2t,\\
E=1360\cdot\pi\cdot({6370\cdot10^3})^2\cdot(3600\cdot24)=1.5\cdot10^{22}\text{ J}."


(b) This is equivalent to the following volume of heating oil:


"V=\frac{E}{u}=\frac{1.5\cdot10^{22}}{37.3\cdot10^6}=4.02\cdot10^{14}\text{ l}."

(c) The energy that the Earth receives is


"E_E=E_0\pi R_E^2(1-\alpha)=1.21\cdot10^{17}\text{ W}."

Apply Stefan-Boltzman law to find the temperature:


"E_E=\sigma AT^4=\sigma(4\pi R_E^2)T^4,\\\space\\
T=\sqrt[4]{\frac{E_E}{4\sigma\pi R^2_E}}=254\text{ K, or -19°C}."

(d) As we know, the average temperature on the Earth is 15°C. The difference is caused by the fact that earth's nucleus provides some heat to the planet as well as activity of living organisms. Also, the value of "\alpha", which was taken as 0.3, is not a constant value and depends on weather and other factors.